The equation of motion of a cyclist has the form: x = 40 - 8t. Determine the direction and speed at which the cyclist is moving: *


1
Expert's answer
2021-04-16T07:22:45-0400

To be given in question

x=40 - 8t"\longrightarrow (1)"

To be asked in question

Direction and speed =?

equation (1) take differenciate with respect t

"\frac{dx}{dt}=-8 \longrightarrow (2)"

Speed always positive

"\ |\frac{dx}{dt}|=|-8|"

"Speed (v)=8 meter\/sec"

Speed = 8 meter/sec


Speed direction:-

"\frac{dx}{dt}=-8"

velocity is negetive mean speed direction is negative x-axis

speed direction is negative (-x) direction.